Rene has R60000 invested among a checking account, a savings account and a bond paying, respectively, 2%, 5% and 7% annual interest rate. She invested R23000 less in her savings account than in her bond. She earns a total of R3550 in annual interest. How much did she invest in each account
1
Expert's answer
2020-06-22T17:35:46-0400

let x bonds invested


make up

the equation


0.05x+0.07(x+23000)=((3500-(0.05x+0.07(x+23000))

0.05x+0.07x+1610=((3550-(0.05x+0,07x+1610))

0.12x+1610=((3550-(0.12x+1610))

0.12x+1610=3350-0.12x-1610

0.24x=330

x=1375


bond paying: 1375+23000=24375

savings account:1375

checking account=60 000-1375-24375=34250